OpenStudy (anonymous):

What is the slope and point of y=3x-13?

OpenStudy (anonymous):

In y = mx + b form, the slope will be the number in the m position. Can you tell me the slope now ?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

correct...so you have a slope of 3 as for your points....what exactly do you need ? The x and y intercepts ?

OpenStudy (anonymous):

is it 5,2?

OpenStudy (anonymous):

in y = mx + b, the b is your y intercept....so your y intercept is -13 or (0,-13) to find the x intercept, sub in 0 for y and solve for x y = 3x - 13 0 = 3x - 13 -3x = -13 x = 13/3....so your x intercept is (13/3,0) or if you need any points....no particular ones..(5,2) y = 3x - 13. 2 = 3(5) - 13 2 = 15 - 13 2 = 2 (correct).....(5,2) is also a set of points on the line

OpenStudy (anonymous):

Can you help me with 3 others also? Please and thank you

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

What is the slope and point of y=3x+11?

OpenStudy (anonymous):

i think the slope would be 1/3, right?

OpenStudy (anonymous):

no.... y = mx + b y = 3x + 11 what number is in the m position....that is your slope

OpenStudy (anonymous):

3!

OpenStudy (anonymous):

yep.....and if you want to find a certain point...you can just use the y intercept, which is in the b position. y = mx + b y = 3x + 11 so you y int = (0,11)

OpenStudy (anonymous):

Then the point would be 2,17

OpenStudy (anonymous):

Next question: What is the slope and point of 3y=x+11?

OpenStudy (anonymous):

lets see if that works y = 3x + 11 17 = 3(2) + 11 17 = 6 + 11 17 = 17 (correct) yes...that works as well....(2,17) is also on the line

OpenStudy (anonymous):

3y = x + 11.....lets put this in y = mx + b form divide both sides by 3 y = 1/3x + 11/3 so your slope (number in the m position) is what ?

OpenStudy (anonymous):

1 over 3?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

I don't know the point on this one

OpenStudy (anonymous):

to find any points....just sub ANY number in for x and solve for y y = 1/3x + 11/3......lets make x = 1 y = 1/3(1) + 11/3 y = 1/3 + 11/3 y = 12/3 y = 4 so when x = 1, y = 4...or (1,4) you can use any number to sub in to fins the other number

OpenStudy (anonymous):

typo...you can use any number to sub in to find the other number

OpenStudy (anonymous):

So is the point -5,2?

OpenStudy (anonymous):

lets see if that works.. y = 1/3x + 11/3 2 = 1/3(-5) + 11/3 2 = - 5/3 + 11/3 2 = 6/3 2 = 2 (correct) (-5,2) ...that works as well

OpenStudy (anonymous):

Last question: What is the point and slope of 3y=x-13? The slope is 1 over 3

OpenStudy (anonymous):

3y = x - 13 y = 1/3x - 13/3 yep...slope is 1/3 can you find a point on the line ?

OpenStudy (anonymous):

you can sub in ANY number for y and solve for x...or you can sub in ANY number for x and solve for y.....or you can just use the y int...which is the number in the b position...there are several different points on this line

OpenStudy (anonymous):

the y int (the number in the b position)....(0,-13/3) is a point on this line

OpenStudy (anonymous):

anymore questions ?

OpenStudy (anonymous):

thank you for the testimonial :)